You purchased one share of Kohl's Corporation, Inc for $55.64 per share. The company paid a dividend of $3.37 per share during the year, and had an ending share price of $54.31. What is the capital gains yield?
Answer #1
Capital gains Yield (CGY)= (Ending price-Purchase price)/Purchase price
ie.(54.31-55.64)/55.64=
-2.39%
(ANSWER)
Dividend yield(DY)= $ dividend/Current stock price
3.37/55.64=
6.06%
Total Holding period yield=
(Ending Price+ $ dividend-Purchase price)/Purchase price
ie. (54.31+3.37-55.64)/55.64=
3.67%
ie. CPGY+DY
-2.39%+6.06%=
3.67%